# Quantum One-Time Programs, Revisited

### v1

@@ 1.1 Our Results
1.2 Concurrent Work and Related Works
Concurrent Work.
One-Time Programs.
Signature Tokens.
Signature tokens are a special case of one-time programs that allow the evaluator to sign a single message, and no more. They were proposed by Ben-David and Sattath [BDS23] in the oracle model and subsequently generalized to the plain model using indistinguishability obfuscation [CLLZ21a]. Both of these works consider a very specific form of one-time security: an adversarial evaluator should not be able to output two (whole) valid signatures.
@@ 2.1 Definitional Works
In the single-effective-query model, we relax $\mathsf{Sim}$’s single-physical-query restriction to allow multiple queries, as long as they are "gentle" or repeated. We will define a stateful oracle $f_{\mathsf{SEQ}}$ which tracks at all times which evaluations $f(x;r)$ the adversary has knowledge about. If $f_{\mathsf{SEQ}}$ receives a query to some $x^{\prime}$ while it knows the adversary has knowledge about an evaluation on $x\neq x^{\prime}$, it will refuse to answer. Using this oracle, we may define single-effective query simulation-security in the same manner as our previous attempt by giving the simulator access to the single-effective-query oracle $f_{\mathsf{SEQ}}$ instead of the single-physical-query oracle $f_{1}$:
@@
|$\ket{x^{*},u}_{\mathcal{Q}}\otimes\ket{H_{\emptyset}}\overset{U_{f_{\$}}}{\longrightarrow}\sum_{r}\ket{x^{*},u\oplus f(x^{*},H(x^{*}))}_{\mathcal{Q}}\otimes\ket{H(x^{*})}_{\mathcal{H}_{x^{*}}}\otimes\sum_{H:H(x^{*})=r}\bigotimes_{x\in\mathcal{X},x\neq x^{*}}\ket{H(x)}_{\mathcal{H}_{x}}$
@@ 2.2 Positive Results
Our construction in the oracle model is inspired by the use of the "hidden subspace states" in the literatures of quantum money [AC12], signature tokens and quantum copy protection [BDS23, CLLZ21b]. A subspace state |$\ket{A}$ is a uniform superposition over all vectors in some randomly chosen, secret subspace $A\subset\mathbb{F}_{2}^{\lambda}$. Specifically, |$\ket{A}\propto\sum_{v\in A}\ket{v}$, where dimension of $A$ is $\lambda/2$ and $\lambda$ is the security parameter. These parameters ensure that $A$ has exponentially many elements but is still exponentially small compared to the entire space.
@@ 4.2 Single effective query (SEQ) simulation-based one-time security
In this section, we describe an alternative model of access which we call the single-effective-query model. This model intuitively allows the adversary to make multiple queries, but it can only have information about one evaluation at a time.
@@
Let $f:\mathcal{X}\times\mathcal{R}\to\mathcal{Y}$ be a randomized function, where the user specifies $x$ and $r$ is chosen randomly. [^6] In general, this function may or may not output its randomness $r$. We define several oracles to handle queries to $f$:
  • $O_{f}$ answers any query $(x,r)$ with $f(x,r)$.
  • $f_{\$}$: on receiving a query $x$, it uses a random oracle $h$ to sample the randomness $r=h(x)$ used for computing $f(x,r)$.
  • $f_{\$,1}$ additionally augments $f_{\$}$ with the ability to recognize how many distinct (effective) queries the user has made. If it receives a second query $x^{\prime}$, the oracle refuses to answer.
The formal definitions of these oracles are below.
Randomized Oracle $f_{\$}$.
Let $f_{\$}$ be an oracle for $f$ that samples a random oracle $h$ and answers each query $x$ with $f(x,h(x))$. The function $h$ ensures that each input $x$ gets a fresh random $r$.
  1. 1.
    The oracle begins with an oracle register $\mathcal{H}$ that holds the following uniform superposition over all random oracles $h:\mathcal{X}\to\mathcal{R}$, represented as a truth table with $|\mathcal{X}|$ entries:
    |$\ket{H_{\emptyset}}:=\frac{1}{|\mathcal{R}|^{|\mathcal{X}|/2}}\sum_{h:\mathcal{X}\to\mathcal{R}}\ket{h}$
  2. 2.
    The user submits a superposition over queries of the form |$\ket{x,u}_{\mathcal{Q}}$, where $x\in\mathcal{X}$ and $u\in\mathcal{Y}$.
  3. 3.
    The oracle applies a unitary $U_{f_{\$}}$ to registers $\mathcal{H}$ and $\mathcal{Q}$ that maps:
    |$\ket{x,u}_{\mathcal{Q}}\otimes\ket{h}_{\mathcal{H}}\overset{U_{f_{\$}}}{\longrightarrow}\ket{x,u\oplus f(x,h(x))}_{\mathcal{Q}}\otimes\ket{h}_{\mathcal{H}}$
We note that the initial superposition |$\ket{H_{\emptyset}}$ can be alternatively represented using Zhandry’s compressed oracle technique [Zha19a].
Recording Queries.
After a single query has been answered, the $\mathcal{Q}$ and $\mathcal{H}$ registers are entangled. If we trace out the $\mathcal{Q}$ register, the state of $\mathcal{H}$ is now a mixed state that reveals which input $x$ was queried.
[…]
\$\mathcal{H}\backslash x=\bigtimes_{x^{\prime}\in\mathcal{X}\backslash\{x\}}\mathcal{H}_{x^{\prime}}$
Next,
Let |$\text{Let }\ket{R}$ $=\frac{1}{\sqrt{|\mathcal{R}|}}\cdot\sum_{r\in\mathcal{R}}\ket{r}$
Then |$\text{Then }\ket{H_{\emptyset}}$ $=\bigotimes_{x\in\mathcal{X}}\ket{R}_{\mathcal{H}_{x}}$
After $f_{\$}$ has responded to a single query |$\ket{x,0}$, the state on the $\mathcal{H}_{x}$ register is the following mixed state $\rho_{x}$:
$\rho_{x}$ $:=\frac{1}{|\mathcal{R}|}\cdot\sum_{r,r^{\prime}\in\mathcal{R}}\ket{r}\bra{r^{\prime}}\cdot\mathbbm{1}_{f(x,r)=f(x,r^{\prime})}$
and the state on the $\mathcal{H}$ register is $\sigma^{H}_{x}$:
$\sigma^{H}_{x}$ $=\bigotimes_{x^{\prime}\in\mathcal{X}\backslash\{x\}}\left(\ket{R}\bra{R}\right)_{\mathcal{H}_{x^{\prime}}}\otimes\left(\rho_{x}\right)_{\mathcal{H}_{x}}$
If $f$ has sufficiently high min-entropy, then $\rho_{x}$ is distinguishable from |$\ket{R}\bra{R}$ | with high advantage (C.1).
Single-Effective-Query Oracle $f_{\$,1}$.
@@
We can define an oracle $f_{\$,1}$ that does not allow the number of recorded queries to be greater than $1$. If answering the query would increase the number of recorded queries to $2$ or more, then $f_{\$,1}$ does not answer it. Also, when $f_{\$,1}$ does answer a query, it will flip a bit $b$ to indicate that it has done so.
  1. 1.
    $f_{\$,1}$ initializes the oracle register $\mathcal{H}$ to |$\ket{H_{\emptyset}}$.
  2. 2.
    The user submits a superposition over states of the form |$\ket{x,u,b}_{\mathcal{Q}}$ in register $\mathcal{Q}=(\mathcal{Q}_{x},\mathcal{Q}_{u},\mathcal{Q}_{b})$, where $x\in\mathcal{X}$, $u\in\mathcal{Y}$, and $b\in\{0,1\}$.
  3. 3.
    $f_{\$,1}$ responds by applying a unitary $U_{f_{\$,1}}$ to the $\mathcal{Q}$ and $\mathcal{H}$ registers, which acts as follows. Let |$\ket{\psi}$ be orthogonal to |$\ket{R}^{\otimes(|\mathcal{X}|-1)}$. Then:
    |$\𝑥\displaystyle\ket{x,u,b}_{\mathcal{Q}}\otimes\ket{r}_{\mathcal{H}_{x}}\otimes\ket{R}^{\otimes(|\mathcal{X}|-1)}_{\mathcal{H}\backslash x}$ $\𝑥\displaystyle\overset{U_{f_{\$,1}}}{\longrightarrow}\ket{x,u\oplus f(x,r),b\oplus 1}_{\mathcal{Q}}\otimes\ket{r}_{\mathcal{H}_{x}}\otimes\ket{R}^{\otimes(|\mathcal{X}|-1)}_{\mathcal{H}\backslash x}$
    |$\𝑥\displaystyle\ket{x,u,b}_{\mathcal{Q}}\otimes\ket{r}_{\mathcal{H}_{x}}\otimes\ket{\psi}_{\mathcal{H}\backslash x}$ $\𝑥\displaystyle\overset{U_{f_{\$,1}}}{\longrightarrow}\ket{x,u,b}_{\mathcal{Q}}\otimes\ket{r}_{\mathcal{H}_{x}}\otimes\ket{\psi}_{\mathcal{H}\backslash x}$
@@
We can also describe $f_{\$,1}$ by implementing the oracle $H$ using the compressed oracle described in Section 3.3, instead of its expanded version. In this case, the register $\mathcal{H}$ contains a superposition over databases $D$, which are sets of input/output pairs $(x,y)$ (see Section 3.3 for more details). $\mathcal{H}$ is initialized to $\ket{\emptyset}$ when $f_{\$,1}$ is initialized.
@@
We remark that these implementations can be seen to be equivalent by inspecting the compressed oracle definition (see Section 3.3). In particular, the oracle states can be mapped to each other by applying the compression operation $\mathsf{Decomp}_{x}$ on every possible input $x$. Consider the expanded oracle state |$\𝑥\ket{r}_{\mathcal{H}_{x}}\otimes\ket{\psi}_{\mathcal{H}\backslash x}$. If |$\ket{\psi}_{\mathcal{H}\backslash x}=\ket{R}^{\otimes(|\mathcal{X}|-1)}$, then the resulting compressed database will not contain any $(x^{\prime},r^{\prime})$ where $x^{\prime}\neq x$, so both versions of $f_{\$,1}$ will evaluate $f$. On the other hand, if |$\𝑥\ket{\psi}_{\mathcal{H}\backslash x}$ is orthogonal to |$\ket{R}^{\otimes(|\mathcal{X}|-1)}$, then the resulting compressed database will contain at least one entry $(x^{\prime},r)$ where $x^{\prime}\neq x$, so both versions will operate as the identity.
Using the compressed oracle representation, it is not too hard to see that $\mathcal{H}$ records at most one query at a time.
@@ 5.3 Which Functions is SEQ Access Meaningful For?
To prove this claim, we introduce the following technical lemma about compositions $f\circ H$ of random functions where $f$ and $H$ are implemented as compressed oracles. It shows that if $f$ records a query $x\|y$, then $H$ must record a corresponding query $x$ where $f(x)=y$. Intuitively, this implies that $f$ can only record a single query, since that is the restriction on $H$. Any input/output pairs that the adversary learns will be recorded by $f$, so they can only learn a single one. We prove the technical lemma in Section D.2.111111Section D.2 also contains a related technical lemma which shows that if an adversary has access to $H\circ G$ and $H$ records an entry $(y,z)$, then $G$ records an entry $(x,y)$ for some $x$. The difference from the lemma mentioned here is that $H$ does not also take $G$’s input $x$ as part of its input.
@@
Let us denote $p:=\mathcal{A}^{f_{\$,1}}(1^{\lambda})\to(x_{1},r_{1},y_{1}=f(x_{1},r_{1})),(x_{2},r_{2},y_{2}=f(x_{2},r_{2}))$. By the above observation, we can also denote $p$ as $\mathcal{A}^{\prime}$’s advantage of outputting $(x_{1},r_{1},y_{1}=f(x_{1},r_{1})),(x_{2},r_{2},y_{2}=f(x_{2},r_{2}))$. By Lemma 3.4, we have $\sqrt{p}\leq\sqrt{p^{\prime}}+\sqrt{2/|\mathcal{Y}|}$, where $p^{\prime}$ is the probability that $D_{f}(x_{1},r_{1})=y_{1},D_{f}(x_{2},r_{2})=y_{2}$ after a computational basis measurement on $D_{f}$. Lemma D.4 shows that whenever $D_{f}(x_{1},r_{1})=y_{1}$ and $D_{f}(x_{2},r_{2})=y_{2}$, there are corresponding entries $D_{H}(x_{1})=r_{1}$ and $D_{H}(x_{2})=r_{2}$ in $H$’s compressed database, except with probability $4q^{2}\left(2/|\mathcal{R}|-1/|\mathcal{R}|^{2}\right)$, where $q$ is the number of queries the adversary has made. Since $D_{H}$ contains at most one entry at a time (4.7) and $\mathcal{R}$ has superpolynomial size, $p^{\prime}$ must be negligible in $\lambda$. Since $\mathcal{Y}$ also has superpolynomial size, $p$ must be negligible as well. ∎
@@
Let =$\mathcal{F}=\{f:\mathcal{X}\times\mathcal{R}\rightarrow\mathcal{Y}\}_{f}$ be a function family associated with distribution $D$. Let $G:\mathcal{X}\times\mathcal{R}^{\prime}\rightarrow\mathcal{R}$ be a random function where $|\mathcal{X}|\cdot|\mathcal{R}^{\prime}|^{2}/|\mathcal{R}|=\mathsf{negl}(\lambda)$ and define $f_{G}:\mathcal{X}\times\mathcal{R}^{\prime}\rightarrow\mathcal{Y}$ by $f_{G}(x;r)=f(x;G(x,r))$. Define $\mathsf{Distr}_{G}$ as the distribution which samples $f\leftarrow D$, then outputs $f_{G}$.
@@
To prove this claim, we use another technical lemma which shows that adversaries who can find images of $G$ must know corresponding preimages. Intuitively, it will allow us to show that an adversary who finds an input/output pair $((x,r),y)$ must know an $r^{\prime}$ such that $G(x,r^{\prime})=r$.
@@
Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be a random function where $|\mathcal{X}|<|\mathcal{Y}|$. Consider an oracle algorithm $A$ makes $q$ of queries to $G$, then outputs a set of $k$ values $\mathbf{y}=(y_{1},\dots,y_{k})$. Let $p$ be the probability that for every $i$, there exists an $x_{i}\in\mathcal{X}$ such that $G(x_{i})=y_{i}$.
Now consider running the same experiment where $G$ is instead implemented as a compressed oracle, and measuring its database register after $A$ outputs to obtain $D$. Let $p^{\prime}$ be the probability that for every $i$, there exists an $x_{i}\in\mathcal{X}$ such that $D(x_{i})=y_{i}$. If $k$ and $q$ are $\mathsf{poly}(\lambda)$ and $|\mathcal{X}|^{k}/|\mathcal{Y}|=\mathsf{negl}(\lambda)$, then [^12] We remark that the reliance on the number of queries is unlikely to be tight. A tighter bound might be achieved by performing a direct computation of the effects of querying $G$ on every $x\in\mathcal{X}$ at the end of the experiment.
@@
We first claim that no adversary can output a tuple $((x,r),y)$ such that $f(x,r)=y$ and $r$ is not in the image of $G(x,\cdot)$, except with negligible probability. This follows from quantum blind unforgeability along with the observation that the SEQ oracle for $f_{G}$ can be evaluated using a list of all evaluations $f(x,G(m,r^{\prime}))$ for $r^{\prime}\in\mathcal{R}^{\prime}$.
Second, we claim that if an adversary outputs two valid input/output tuples $((x_{1},r_{1}),y_{1})$ and $((x_{2},r_{2}),y_{2})$ where $x_{1}\neq x_{2}$, then with overwhelming probability $r_{1}\neq r_{2}$. Say the adversary did so with probability $p$. If $r_{1}=r_{2}$, then by the previous claim, there exist $r_{1}^{\prime}$ and $r_{2}^{\prime}$ such that $G(x_{1},r_{1}^{\prime})=G(x_{2},r_{2}^{\prime})$. Thus, we could find a collision in $G$ with probability $p\cdot|\mathcal{R}^{\prime}|^{2}$ by guessing $r_{1}^{\prime}$ and $r_{2}^{\prime}$. Lemma 3.5 shows that after $q$ queries, the probability of finding a collision is $O(q^{3}/|\mathcal{R}|)$. Since $|\mathcal{R}^{\prime}|^{2}/|\mathcal{R}|=\mathsf{negl}(\lambda)$, we have $p\leq O(q^{3}|\mathcal{R}^{2}|/|\mathcal{R}|)=\mathsf{negl}(\lambda)$.
Lemma D.5 shows that if we were to implement $G$ as a compressed oracle, then whenever the adversary finds two valid input/output tuples where $r_{1}\neq r_{2}$, with overwhelming probability $G$’s compressed database contains entries $(x_{1}^{\prime}\|r_{1}^{\prime},r_{1})$ and $(x_{2}^{\prime}\|r_{2}^{\prime},r_{2})$. By Lemma 5.13, whenever this occurs, $H$ also contains two entries, except with negligible probability. However, $H$ never contains more than one entry (4.7). Combining all of these facts together, any QPT adversary given SEQ access to $f_{G}\leftarrow\mathsf{Distr}_{G}$ cannot output two distinct tuples $((x_{1},r_{1}),y_{1})$ and $((x_{2},r_{2}),y_{2})$ such that $f(x_{1},r_{1})=y_{1}$ and $f(x_{2},r_{2})=y_{2}$, except with negligible probability. ∎

### v2

@@ 2.1 Definitional Works
In the single-effective-query model, we relax $\mathsf{Sim}$’s single-physical-query restriction to also allow multiple queries, as long as they are "gentle" or repeated. We will define a stateful oracle $f_{\mathsf{SEQ}}$ which tracks at all times which evaluations $f(x;r)$ the adversary has knowledge about. If $f_{\mathsf{SEQ}}$ receives a query to some $x^{\prime}$ while it knows the adversary has knowledge about an evaluation on $x\neq x^{\prime}$, it will refuse to answer. Using this oracle, we may define single-effective query simulation-security in the same manner as our previous attempt by giving the simulator access to the single-effective-query oracle $f_{\mathsf{SEQ}}$ instead of the single-physical-query oracle $f_{1}$:
@@
|$\ket{x^{*},u}_{\mathcal{Q}}\otimes\ket{H_{\emptyset}}\overset{U_{f_{\$}}}{\longrightarrow}\sum_{r}\ket{x^{*},u\oplus f(x^{*},r)}_{\mathcal{Q}}\otimes\ket{r}_{\mathcal{H}_{x^{*}}}\otimes\sum_{H:H(x^{*})=r}\bigotimes_{x\in\mathcal{X},x\neq x^{*}}\ket{H(x)}_{\mathcal{H}_{x}}$
@@ 2.2 Positive Results
Our construction in the oracle model is inspired by the use of the "hidden subspace states" in the literatures of quantum money [AC12], signature tokens and quantum copy protection [BDS23, CLLZ21a]. A subspace state |$\ket{A}$ is a uniform superposition over all vectors in some randomly chosen, secret subspace $A\subset\mathbb{F}_{2}^{\lambda}$. Specifically, |$\ket{A}\propto\sum_{v\in A}\ket{v}$, where dimension of $A$ is $\lambda/2$ and $\lambda$ is the security parameter. These parameters ensure that $A$ has exponentially many elements but is still exponentially small compared to the entire space.
@@
More Applications and Implications: Generic Signature Tokens, NIZK, Quantum Money
We show a way to lift signature schemes that satisfy a property called blind unforgeability to possess one-time security. Unlike the [BDS23] signature token scheme where the verification key has to be updated once we delegate a one-time signing token to some delegatee, our signature token scheme can use an existing public verification procedure.
Apart from the above one-time PRF in the plain model, we also instantiate one-time NIZK from iO and LWE in the plain model, using a similar construction as in the one-time PRF and the NIZK from iO in [SW14]. The proof requires more careful handling because the NIZK proof also has publicly-verifiable property.
Finally, we show that one-time program for publicly verifiable programs (e.g. signature, NIZK) implies public-key quantum money. Despite the destructible nature of the one-time program, we can design a public verification procedure that gently tests a program’s capability of computing a function and use a one-time program token state as the banknote.
@@ 2.3 Impossibility Results
2.4 Concurrent Work and Related Works
Concurrent Work.
One-Time Programs.
Signature Tokens.
Signature tokens are a special case of one-time programs that allow the evaluator to sign a single message, and no more. They were proposed by Ben-David and Sattath [BDS23] in the oracle model and subsequently generalized to the plain model using indistinguishability obfuscation [CLLZ21b]. Both of these works consider a very specific form of one-time security: an adversarial evaluator should not be able to output two (whole) valid signatures.
@@ 4.2 Single effective query (SEQ) simulation-based one-time security
In this section, we describe an alternative model of access which we call the single-effective-query (SEQ) model. This model intuitively allows the adversary to make multiple queries, but it can only have information about one evaluation at a time.
For example, the adversary can make a query, then uncompute it, and then make a query on a different input. After any query, the adversary only has information about at most one point of the function, so they never make more than a single effective query. We will use the compressed oracle technique to record queries the adversary makes and enforce that they cannot make more than a single effective query.
@@
Let $f:\mathcal{X}\times\mathcal{R}\to\mathcal{Y}$ be a randomized function, where the user specifies $x$ and $r$ is chosen randomly [^6] In general, this function may or may not output its randomness $r$.. Let $f_{\$}$ be an implementation of $f$ that uses a random oracle $H$ to compute the randomness. On receiving input $x$, $f_{\$}$ computes $r=H(x)$ and outputs
$f_{\$}(x)=f(x,H(x))$
@@
The single effective query (SEQ) oracle $f_{\$,1}$ augments $f_{\$}$ with the ability to recognize how many distinct (effective) queries the user has made to $H$ by implementing $H$ as a compressed oracle that records the user’s queries. If answering a given query would increase the number of recorded queries to $2$ or more, then $f_{\$,1}$ does not answer it. Also, when $f_{\$,1}$ does answer a query, it will flip a bit $b$ to indicate that it has done so. The formal definition $f_{\$,1}$ is given below.
@@
We define $f_{\$,1}$ to implement the random oracle $H$ using the compressed oracle technique described in Section 3.3. In this case, the oracle’s $\mathcal{H}$ register contains a superposition over databases $D$, which are sets of input/output pairs $(x,r)$. When $f_{\$,1}$ is initialized, $\mathcal{H}$ is initialized to $\ket{\emptyset}$.
@@
It is not hard to see that $\mathcal{H}$ records at most one query at a time.
@@ 5.3 Which Functions is SEQ Access Meaningful For?
To prove this claim, we introduce the following technical lemma about compositions $f\circ H$ of random functions where $f$ and $H$ are implemented as compressed oracles. It shows that if $f$ records a query $x\|y$, then $H$ must record a corresponding query $x$ where $f(x)=y$. Intuitively, this implies that $f$ can only record a single query, since that is the restriction on $H$. Any input/output pairs that the adversary learns will be recorded by $f$, so they can only learn a single one. We prove the technical lemma in Section C.2.111111Section C.2 also contains a related technical lemma which shows that if an adversary has access to $H\circ G$ and $H$ records an entry $(y,z)$, then $G$ records an entry $(x,y)$ for some $x$. The difference from the lemma mentioned here is that $H$ does not also take $G$’s input $x$ as part of its input.
@@
Let us denote $p:=\mathcal{A}^{f_{\$,1}}(1^{\lambda})\to(x_{1},r_{1},y_{1}=f(x_{1},r_{1})),(x_{2},r_{2},y_{2}=f(x_{2},r_{2}))$. By the above observation, we can also denote $p$ as $\mathcal{A}^{\prime}$’s advantage of outputting $(x_{1},r_{1},y_{1}=f(x_{1},r_{1})),(x_{2},r_{2},y_{2}=f(x_{2},r_{2}))$. By Lemma 3.4, we have $\sqrt{p}\leq\sqrt{p^{\prime}}+\sqrt{2/|\mathcal{Y}|}$, where $p^{\prime}$ is the probability that $D_{f}(x_{1},r_{1})=y_{1},D_{f}(x_{2},r_{2})=y_{2}$ after a computational basis measurement on $D_{f}$. Lemma C.4 shows that whenever $D_{f}(x_{1},r_{1})=y_{1}$ and $D_{f}(x_{2},r_{2})=y_{2}$, there are corresponding entries $D_{H}(x_{1})=r_{1}$ and $D_{H}(x_{2})=r_{2}$ in $H$’s compressed database, except with probability $4q^{2}\left(2/|\mathcal{R}|-1/|\mathcal{R}|^{2}\right)$, where $q$ is the number of queries the adversary has made. Since $D_{H}$ contains at most one entry at a time (4.7) and $\mathcal{R}$ has superpolynomial size, $p^{\prime}$ must be negligible in $\lambda$. Since $\mathcal{Y}$ also has superpolynomial size, $p$ must be negligible as well. ∎
@@
Let =$\mathcal{F}=\{f:\mathcal{X}\times\mathcal{R}\rightarrow\mathcal{Y}\}_{f}$ be a function family associated with distribution $D$. Let $G:\mathcal{X}\times\mathcal{R}^{\prime}\rightarrow\mathcal{R}$ be a random function where $|\mathcal{R}^{\prime}|^{2}/|\mathcal{R}|=\mathsf{negl}(\lambda)$ and define $f_{G}:\mathcal{X}\times\mathcal{R}^{\prime}\rightarrow\mathcal{Y}$ by $f_{G}(x;r)=f(x;G(x,r))$. Define $\mathsf{Distr}_{G}$ as the distribution which samples $f\leftarrow D$, then outputs $f_{G}$.
@@
To prove this claim, we use another technical lemma which shows that adversaries who can find images of $G$ must know corresponding preimages. Intuitively, it will allow us to show that an adversary who finds an input/output pair $((x,r),y)$ must know an $r^{\prime}$ such that $G(x,r^{\prime})=r$. We prove the technical lemma in Section C.3.
@@
Let $G:\mathcal{X}_{1}\times\mathcal{X}_{2}\rightarrow\mathcal{Y}$ be a random function where $|\mathcal{X}_{2}|<|\mathcal{Y}|$. Consider an oracle algorithm $A$ makes $q$ of queries to $G$, then outputs two vector of $k$ values $\mathbf{x^{(1)}}=\left(x_{1}^{(1)}\dots,x_{k}^{(1)}\right)$ and $\mathbf{y}=(y_{1},\dots,y_{k})$. Let $p$ be the probability that for every $i$, there exists an $x_{i}^{(2)}\in\mathcal{X}$ such that $G\left(x_{i}^{(1)},x_{i}^{(2)}\right)=y_{i}$.
Now consider running the same experiment where $G$ is instead implemented as a compressed oracle, and measuring its database register after $A$ outputs to obtain $D$. Let $p^{\prime}$ be the probability that for every $i$, there exists an $x_{i}^{(2)}\in\mathcal{X}_{2}$ such that $D\left(x_{i}^{(1)}\|x_{i}^{(2)}\right)=y_{i}$. If $k$ and $q$ are $\mathsf{poly}(\lambda)$ and $|\mathcal{X}_{2}|^{k}/|\mathcal{Y}|=\mathsf{negl}(\lambda)$, then [^12] We remark that the reliance on the number of queries is unlikely to be tight. However, this bound is sufficient for our purposes since we will anyway combine it with results that require a query-bounded adversary.
@@
We first claim that no adversary can output a tuple $((x,r),y)$ such that $f(x,r)=y$ and $r$ is not in the image of $G(x,\cdot)$, except with negligible probability. This follows from quantum blind unforgeability along with the observation that the SEQ oracle for $f_{G}$ can be evaluated using a list of all evaluations $f(x,G(x,r^{\prime}))$ for $r^{\prime}\in\mathcal{R}^{\prime}$.
Second, we claim that if an adversary outputs two valid input/output tuples $((x_{1},r_{1}),y_{1})$ and $((x_{2},r_{2}),y_{2})$ where $x_{1}\neq x_{2}$, then with overwhelming probability $r_{1}\neq r_{2}$. Say the adversary did so with probability $p$. By the previous claim, whenever the adversary succeeds, there exist $r_{1}^{\prime}$ and $r_{2}^{\prime}$ such that $r_{1}=G(x_{1},r_{1}^{\prime})$ and $r_{2}=G(x_{2},r_{2}^{\prime})$, except with negligible probability. If $r_{1}=r_{2}$, then $G(x_{1},r_{1}^{\prime})=G(x_{2},r_{2}^{\prime})$. Thus, we could find a collision in $G$ with probability $p/|\mathcal{R}^{\prime}|^{2}-\mathsf{negl}(\lambda)$ by guessing $r_{1}^{\prime}$ and $r_{2}^{\prime}$. Lemma 3.5 shows that after $q$ queries, the probability of finding a collision is $O(q^{3}/|\mathcal{R}|)$. Since $|\mathcal{R}^{\prime}|^{2}/|\mathcal{R}|=\mathsf{negl}(\lambda)$, we have $p\leq O(q^{3}|\mathcal{R}^{2}|/|\mathcal{R}|)=\mathsf{negl}(\lambda)$.
Lemma 5.17 shows that if we were to implement $G$ as a compressed oracle, then whenever the adversary finds two valid input/output tuples where $r_{1}\neq r_{2}$, with overwhelming probability $G$’s compressed database contains entries $(x_{1}^{\prime}\|r_{1}^{\prime},r_{1})$ and $(x_{2}^{\prime}\|r_{2}^{\prime},r_{2})$. By Lemma 5.13, whenever this occurs, $H$ also contains two entries, except with negligible probability. However, $H$ never contains more than one entry (4.7). Combining all of these facts together, any QPT adversary given SEQ access to $f_{G}\leftarrow\mathsf{Distr}_{G}$ cannot output two distinct tuples $((x_{1},r_{1}),y_{1})$ and $((x_{2},r_{2}),y_{2})$ such that $f(x_{1},r_{1})=y_{1}$ and $f(x_{2},r_{2})=y_{2}$, except with negligible probability. ∎